# No Compromise in Solution Quality: Speeding Up Belief-dependent Continuous POMDPs via Adaptive Multilevel Simplification

### v1

@@ 1.3 Paper Organization
The remainder of this paper is structured as follows. Section 2 provides background in terms of POMDPs, theoretical objective and commonly used objective estimators. We devote Section 3 to our general adaptive multi-level simplification framework. In Section 4 we consider a given belief tree setting in which the belief tree construction is not coupled with the solution. In Section 5 we delve into the MCTS approach in the context of our multilevel simplification. In Section 6 we consider a specific simplification and develop novel bounds on an information-theoretic reward function. Finally, Section 7 presents simulations and results corroborating our ideas. In order not to disrupt the flow of the presentation, proofs are presented in appropriate Appendices.
@@ 2 Background
In this section we present the background.
@@ A POMDP is a tuple
In our generalized formulation, the reward is a function of two subsequent beliefs, an action and an observation:
@@ 2.2 Theoretical Objective
The decision making goal is to find an optimal policy $\pi_{k+}$ maximizing
$\begin{gathered}\!\!\!\!V\!(b_{k},\!\pi_{k+})\!=\!\!\!\!\!\!\!\!\!\underset{z_{k+1:k+L}}{\mathbb{E}}\!\Big{[}\sum_{\ell=k}^{k+L-1}\!\!\!\!\!\gamma^{\ell-k}\!\rho(b_{\ell},\!\pi_{\ell}(b_{\ell}),\!z_{\ell+1},\!b_{\ell+1})|b_{k},\!\pi_{k+}\!\Big{]}\!\!\!\!\\ \text{s.t.}\ \ b_{\ell+1}=\psi(b_{\ell},\pi_{\ell}(b_{\ell}),z_{{\ell}+1}),\end{gathered}$ (6)
where $\psi$ is the Bayesian belief update method. Utilizing the Bellman formulation (6) takes the form of
@@ 2.3.1 Objective Estimator in Case of a Given Belief Tree
$\begin{gathered}\hat{V}(b_{k},\pi_{k+})\!=\!\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\rho(b_{k},\pi_{k}(b_{k}),z^{i}_{k+1},b^{i}_{k+1})+\\ +\gamma\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}V(\psi(b_{k},\pi_{k}(b_{k}),z^{i}_{k+1}),\pi_{(k+1)+}),\end{gathered}$ (13)
and the corresponding estimator for (8) under optimal future policy reads
$\begin{gathered}\hat{Q}(b_{k},a_{k})=\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\rho(b_{k},a_{k},z^{i}_{k+1},b^{i}_{k+1})+\\ +\gamma\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}V(\psi(b_{k},a_{k},z^{i}_{k+1}),\pi^{*}_{(k+1)+}),\end{gathered}$ (14)
@@ 2.3.2 Coupled Action-Value Function Estimation and Belief Tree Construction
where $N(h)$ is the visitation count of the belief node defined by history $h$, $N(ha)$ is the visitation count of the belief-action node, $c$ is the exploration parameter and, $\hat{Q}(ha)$ is the estimator of the action-value function $Q$ for node $ha$ obtained by simulations. When the action is selected, a question arises either to open a new branch in terms of observation and posterior belief or to continue through one of the existing branches. In continuous action, and observation spaces, this can be resolved by the Double Progressive Widening (DPW) technique (Sunberg and Kochenderfer, 2018; Auger et al., 2013). If a new branch is expanded, an observation $o$ is created from state $x$ drawn from the belief $b$.
Let the return, corresponding to lace $i$ starting from some belief $b^{i}_{\ell}$ at depth $\ell-k$, be $g(b^{i}_{\ell})$ for $\ell\in[k:k+L-1]$. More specifically, suppose the new posterior belief was expanded at depth $d^{i}$ of the belief tree such that $d^{i}>\ell$. We have that
$g(b^{i}_{\ell},a_{\ell})=$
$=\underbrace{\rho(b^{i}_{\ell},a_{\ell},b^{i}_{\ell+1})+\sum_{l=\ell+1}^{k+d^{i}}\gamma^{l-\ell}\rho(b^{i}_{l},\pi^{*,i}_{l}(b_{l}),b^{i}_{l+1})}_{\text{belief tree}}+$ (16)
@@ 3.1 Theoretical Simplification Formulation
Simplification is any kind of relaxation of POMDP tuple (1) elements, accompanied by guarantees that quantify the (worst-case or potential) impact of a particular simplification technique on planning performance. In this paper, we consider a specific manifestation of this general simplification framework, which will become clear shortly.
@@
Henceforth in order to avoid unnecessary clutter we will omit the dependence on the observation and denote the bounds over the reward using simplified beliefs as follows
@@
It should be stressed that since in the belief tree $b^{\prime}$ always has a single parent $b$, the reader should think about such a reward as one corresponding to $b^{\prime}$. Since the reward depends on a pair of consecutive beliefs, the same belief is simplified differently for successive rewards $\rho(b,a,b^{\prime})$ and $\rho(b^{\prime},a^{\prime},b^{\prime\prime})$.
@@
Suppose inequality (21) holds for any possible pair of consecutive beliefs, e.g. these are analytical bounds, as opposed to (Zhitnikov and Indelman, 2022). A direct consequence of this fact, alongside the structure of (6), is that
@@
At this point, we want to recall that commonly, the state-dependent element (2) is much easier to calculate than the belief dependent one. Leveraging the structure manifested by (10) the immediate bounds over (3) induce bounds over $Q^{I}(\cdot)$
@@ 3.2 Bounds over the Estimated Objective
Allow us to linger on the $\lambda$ from eq. 10 and 12. It is hard to predict how the objective will behave with various values of $\lambda$. Nevertheless, if the bounds are over the belief-dependent element of the reward, by subtracting (34) from (33), we arrive at
@@
The width of the bounds is monotonically increasing with $\lambda$. Of course, it will also happen to a theoretical analog of such a bounds displayed by eq. (26) and (27). We can envision more speedup with lower values of $\lambda$ and will see it in the simulations.
@@ 3.3 Multi-Level Simplification
If the bounds over the reward are monotonic (assumption 1) and convergent (assumption 2), for both estimators (31) and (32), the bounds on the sample approximation (29) are monotonic as a function of the number of resimplifications and convergent after at most $n_{{\mathrm{max}}}\cdot M$ resimplifciations for any resimplification strategy. Here $M$ is the number of posterior beliefs in (31) or (32) . Namely, if all the rewards are at the maximal simplifciation level $n_{\mathrm{max}}$ we have to reach
@@ 3.4 Adaptive Simplification Mechanics
$j^{\dagger}(b_{\ell})=\operatorname*{arg\,max}_{j}\Big{\{}\underline{\hat{Q}}(b_{\ell},\{a^{j}_{\ell},\pi^{\ast}_{({\ell}+1)+}\})+c^{j}\Big{\}},$ (37)
@@
If the condition displayed by equation (38) is not fullfilled, as depicted in Fig. 3, we shall tighten the bounds (29) by calling a resimplification strategy . Importantly, in case of a given belief tree, even if an overlap is present similar to branch-and-bound technique (Kochenderfer et al., 2022) we can prune any subtree obtained with action $j$ satisfying
@@
We illustrated this aspect in Fig. 3. If the belief tree is constructed gradually as in MCTS based methods and anytime setting, instead of pruning, we still can use (39) to dismiss suboptimal, at current simulation of MCTS, actions.
@@ 3.5 Specific Resimplification Strategies
Simplification level: The resimplification strategy can be directly tied to the simplification level. In this situation the resimplifcation strategy promotes simplification level of the rewards inside the belief tree corresponding to bounds in (28) or (29) based on the simplification level itself. We provide further details in the setting of a given belief tree, considering a PT variant in Section 4].
Gap: Another possibility is that the resimplifciation is tied to the gap $\overline{\rho}^{s}-\underline{\rho}^{s}$. Such a resimplification promotes the simplification level if the reward bounds gap satisfies a certain condition. We describe thoroughly this resimplification flavor in the setting of a given belief tree, considering LAZY variant in Section 4.2, and in MCTS setting, considering a PT variant in Section 5.4.
@@ 4 Adaptive Simplification in the Setting of a Given Belief Tree
The second strategy is described in Section 4.2. This resimplification strategy is based on Gap and it is LAZY strategy. Here, the general idea is to first substitute all the rewards in a given belief tree by bounds with the coarsest simplification level. We then eliminate an overlap between candidate actions only at the root belief node $b_{k}$ by a repetitive descending to the belief tree, promoting the simplification levels along a single lace chosen according to largest gap and ascending back. We emphasize that in this setting, the action-value bounds should not overlap only at the root node in the given belief tree.
Figure 4: Pruning the subtrees by adaptively promoting the simplification levels of the rewards inside. Here the simplification levels of a subtrees are not equal. It is possible that $s^{i}\neq s^{i+1}$. Note that here the superscipts are relative to $b_{\ell}$ as opposed to Fig. 1 and Fig. 5.
As mentioned in the beginning of section 2.3.1, only for simplicity we consider a symmetric setting in terms of sampled actions and the observations but the approach is applicable without any limitations to any given belief tree.
@@ 4.1 Resimplification strategy: PT-SL
We now discuss how the simplification level is updated recursively from the simplifciation level of pertaining reward bounds, and revisit the process to calculate the optimal policy and the corresponding bounds. For some belief node $b_{\ell}$ in the belief tree, consider the bounds $\overline{\hat{Q}}^{s^{j}}(b_{\ell},\{a^{j}_{\ell},\pi^{\ast}_{(\ell+1)+}\})$ and $\underline{\hat{Q}}^{s^{j}}(b_{\ell},\{a^{j}_{\ell},\pi^{\ast}_{(\ell+1)+}\})$ from (44) for different actions $a^{j}_{\ell}\in\mathcal{A}$, that partially overlap and therefore could not be pruned. Each policy tree corresponding to action $a^{j}_{\ell}$ can generally have its own simplification level $s^{j}$. We now iteratively increase the simplification level by $1$. This can be done for each of the branches, if $s^{j}$ is identical for all branches, or only for the branch with the coarsest simplification level.
@@
For the former case, i.e. $s^{i}=s^{j}$, we now have to adapt the simplification level of child tree $i$ to $s^{j}+1$ by calculating the bounds $\overline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$. Here, our key insight is that, instead of calculating these bounds from scratch, we can re-use calculations between different simplification levels, in this case, from level $s^{i}$. As the bounds from that level are available, we can identify only the incremental part that is “missing” to get from simplification level $s^{i}$ to $s^{i}+1$, and update the existing bounds $\overline{\hat{V}}^{s^{i}}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$ to recover $\overline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$ exactly. The same argument applies also for bounds over momentary rewards. In Section 6.2.3 we apply this approach to a specific simplification and reward function.

### v2

@@ 1.3 Paper Organization
The remainder of this paper is structured as follows. Section 2 provides background in terms of POMDPs, theoretical objective and commonly used objective estimators. We devote Section 3 to our general adaptive multi-level simplification framework. In Section 4 we consider a given belief tree setting in which the belief tree construction is not coupled with the solution. In Section 5 we delve into the MCTS approach in the context of our multilevel simplification. In Section 6 we consider a specific simplification and develop novel bounds on an information-theoretic reward function. Section 7 assesses the general adaptation overhead of our methodology. Finally, Section 8 presents simulations and results corroborating our ideas. In order not to disrupt the flow of the presentation, proofs are presented in appropriate Appendices.
@@ 2 Background
In this section we present the background. To elaborate, we present a POMDP with belief dependent rewards followed by theoretical and estimated objectives that correspond to different online POMDP solvers. Our techniques work with estimated objectives.
@@ A POMDP is a tuple
In our generalized formulation, the reward is a function of two subsequent in time beliefs, an action and an observation:
@@ 2.2 Theoretical Objective
The decision making goal is to find an optimal policy $\pi_{k+}$ maximizing the value function as such:
$\begin{gathered}V(b_{k},\pi_{k+})\ \text{s.t.}\ \ b_{\ell+1}=\psi(b_{\ell},\pi_{\ell}(b_{\ell}),z_{{\ell}+1}),\end{gathered}$ (6)
where $V(b_{k},\pi_{k})$ is defined by
$\begin{gathered}\underset{z_{k+1:k+L}}{\mathbb{E}}\!\Big{[}\sum_{\ell=k}^{k+L-1}\!\!\!\!\!\gamma^{\ell-k}\!\rho(b_{\ell},\!\pi_{\ell}(b_{\ell}),\!z_{\ell+1},\!b_{\ell+1})|b_{k},\!\pi_{k+}\!\Big{]}\!\!\!\!\end{gathered}$ (7)
and $\psi$ is the Bayesian belief update method. Utilizing the Bellman formulation (7) takes the form of
@@ 2.3.1 Objective Estimator in Case of a Given Belief Tree
$\begin{gathered}\hat{V}(b_{k},\pi_{k+})\!=\!\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\rho(b_{k},\pi_{k}(b_{k}),z^{i}_{k+1},b^{i}_{k+1})+\\ +\gamma\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\hat{V}(\psi(b_{k},\pi_{k}(b_{k}),z^{i}_{k+1}),\pi_{(k+1)+}),\end{gathered}$ (14)
and the corresponding estimator for (9) under an optimal future policy reads
$\begin{gathered}\hat{Q}(b_{k},a_{k})=\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\rho(b_{k},a_{k},z^{i}_{k+1},b^{i}_{k+1})+\\ +\gamma\frac{1}{n_{z}}\sum_{i=1}^{n_{z}}\hat{V}(\psi(b_{k},a_{k},z^{i}_{k+1}),\pi^{*}_{(k+1)+}),\end{gathered}$ (15)
@@
2.3.2 Interchangeability Between the history and Belief
The purpose of this section is to clarify why further we will use interchangeably belief and the history. The belief is merely a reinterpretation of the knowledge about the POMDP state stored in history in the form of a PDF. The belief $b_{k}$ is a function of the history $h_{k}$. Therefore different histories may yield the same belief. To avoid ambiguity and relate the objectives and their position in the belief tree with some abuse of notation we sometimes switch the dependence on the belief to dependence on corresponding history. In general we can write $b_{\ell}(h_{\ell})$.
@@ 2.3.3 Coupled Action-Value Function Estimation and Belief Tree Construction
where $N(h)$ is the visitation count of the belief node defined by history $h$, $N(ha)$ is the visitation count of the belief-action node, $c$ is the exploration parameter and, $\hat{Q}(ha)$ is the estimator of the action-value function $Q$ for node $ha$ obtained by simulations. The rule described by (16) is a result of modelling exploration as Multi Armed Bandit (MAB) problem (Kocsis and Szepesvári, 2006; Munos, 2014; Auger et al., 2013). When the action is selected, a question arises either to open a new branch in terms of observation and posterior belief or to continue through one of the existing branches. In continuous action, and observation spaces, this can be resolved by the Double Progressive Widening (DPW) technique (Sunberg and Kochenderfer, 2018; Auger et al., 2013). If a new branch is expanded, an observation $o$ is created from state $x$ drawn from the belief $b$.
Let the return, corresponding to lace $i$ starting from some belief $b^{i}_{\ell}$ at depth $\ell-k$, be $g(b^{i}_{\ell},a_{\ell})$ for $\ell\in[k:k+L-1]$. More specifically, suppose the new posterior belief was expanded at depth $d^{i}$ of the belief tree such that $d^{i}>\ell$. We have that $g(b^{i}_{\ell},a_{\ell})$ is composed from two parts, the already expanded tree part and the rollout added part such that
$\!\!\!g(b^{i}_{\ell},a_{\ell}){=}\!$ $\underbrace{\rho(b^{i}_{\ell},a_{\ell},b^{i}_{\ell+1}){+}\!\!\!\!\!\sum_{l=\ell+1}^{k+d^{i}}\!\!\!\!\gamma^{l-\ell}\rho(b^{i}_{l},\pi^{*,i}_{l}(b_{l}),b^{i}_{l+1})}_{\text{belief tree}}\!{+}\!\!\!$ (17)
@@ 3.1 Theoretical Simplification Formulation
Simplification is any kind of relaxation of POMDP tuple (1) elements, accompanied by guarantees that quantify the (worst-case or potential) impact of a particular simplification technique on planning performance. In this section, we present a general simplification framework that is applicable to any reward bounds that satisfy the assumptions stated in Section 3.3.
Our framework applies without any change to parametric and non-parametric beliefs, and to closed-form belief-dependent rewards (that can be calculated exactly, i.e. analytically), as well as to estimated rewards. Therefore, in this paper we do not differentiate between these cases and denote the belief-dependent reward by $\rho(b_{\ell},a_{\ell},z_{\ell+1},b_{\ell+1})$, without using the notation $\hat{\square}$ for estimators. In other words, depending on the setting, $\rho()$ and $b_{\ell}$ can represent, respectively, an analytical reward and a parametric belief, or a reward estimator and a nonparametric belief. In all cases, if one can provide monotonically adaptive bounds on the reward, our framework will return an identical solution as if the decision making was performed with original reward calculations (i.e. depending on the setting, either an analytical reward calculation or reward estimator calculation). In Section 6 we provide a specific incarnation of the framework considering non-parametric beliefs represented by a set of weighted samples and a reward estimator, and where the simplification corresponds to utilizing only a subset of the samples.
@@
Henceforth, in order to avoid unnecessary clutter we will omit the dependence on the observation and denote the bounds over the reward using simplified beliefs as follows
@@
It should be stressed that since in the belief tree $b^{\prime}$ always has a single parent $b$, the reader should think about such a reward as one corresponding to $b^{\prime}$.
@@
Suppose inequality (22) holds for any possible pair of consecutive beliefs, e.g. these are analytical bounds, as opposed to (Zhitnikov and Indelman, 2022b). A direct consequence of this fact, alongside the structure of (7), is that
@@
At this point, we want to recall that commonly, the state-dependent element (2) is much easier to calculate than the belief dependent one. Leveraging the structure manifested by (11) the immediate bounds over (3) induce bounds over $Q^{I}(\cdot)$ as such
@@ 3.2 Bounds over the Estimated Objective
Allow us to linger on the $\lambda$ from eq. 11 and 13. It is hard to predict how the objective will behave with various values of $\lambda$. Nevertheless, if the bounds are over the belief-dependent element of the reward, by subtracting (35) from (34), we arrive at
@@
The width of the bounds is monotonically increasing with $\lambda$. Of course, it will also happen to a theoretical analog of such a bounds displayed by eq. (27) and (28). We can envision more speedup from applying the simplification paradigm with lower values of $\lambda$ and will see it in the simulations.
@@ 3.3 Multi-Level Simplification
If the bounds over the reward are monotonic (assumption 1) and convergent (assumption 2), for both estimators (32) and (33), the bounds on the sample approximation (30) are monotonic as a function of the number of resimplifications and convergent after at most $n_{{\mathrm{max}}}\cdot M$ resimplifications for any resimplification strategy. Here $M$ is the number of posterior beliefs in (32) or (33) . Namely, if all the rewards are at the maximal simplifciation level $n_{\mathrm{max}}$ we have to reach
@@ 3.4 Adaptive Simplification Mechanics
$\!j^{\dagger}(b_{\ell}(h_{\ell})){=}\operatorname*{arg\,max}_{j}\Big{\{}\underline{\!\hat{Q}}(b_{\ell}(h_{\ell}),\!\{a^{j}_{\ell},\pi^{\ast}_{({\ell}+1)+}\}){+}c^{j}(h_{\ell}a^{j})\Big{\}},$ (38)
@@
If the condition displayed by equation (39) is not fulfilled, as depicted in Fig. 3, we shall tighten the bounds (30) by calling a resimplification strategy . Importantly, in case of a given belief tree, even if an overlap is present similar to branch-and-bound technique (Kochenderfer et al., 2022) we can prune any subtree obtained with action $j$ satisfying
@@
We illustrated this aspect in Fig. 3. If the belief tree is constructed gradually as in MCTS based methods and anytime setting, instead of pruning, we still can use (40) to dismiss suboptimal, at current simulation of MCTS, actions (See Fig. 4).
Figure 4: Demonstration of our approach in the setting of MCTS. In contrast to Fig. 3, we cannot prune action $a^{2}$ and can only dismiss it to not participate in resimplifications. This is because, in the next tree queries, $a^{2}$ may be the best action for the robot to take.
@@ 3.5 Specific Resimplification Strategies
Simplification level: The resimplification strategy can be directly tied to the simplification level. In this situation the resimplifcation strategy promotes simplification level of the rewards inside the belief tree corresponding to bounds in (29) or (30) based on the simplification level itself. We provide further details in the setting of a given belief tree, considering a PT variant in Section 4.
Gap: Another possibility is that the resimplification is tied to the gap $\overline{\rho}^{s}-\underline{\rho}^{s}$. Such a resimplification promotes the simplification level if the reward bounds gap satisfies a certain condition. We describe thoroughly this resimplification flavor in the setting of a given belief tree, considering LAZY variant in Section 4.2, and in MCTS setting, considering a PT variant in Section 5.4.
@@ 4 Adaptive Simplification in the Setting of a Given Belief Tree
The second strategy is described in Section 4.2. This resimplification strategy is based on Gap and it is a LAZY strategy. Here, the general idea is to first substitute all the rewards in a given belief tree by bounds with the coarsest simplification level. We then eliminate an overlap between candidate actions only at the root belief node $b_{k}$ by a repetitive descending to the belief tree, promoting the simplification levels along a single lace chosen according to largest gap and ascending back. We emphasize that in this setting, the action-value bounds should not overlap only at the root node in the given belief tree.
Figure 5: Pruning the subtrees by adaptively promoting the simplification levels of the rewards inside. Here the simplification levels of a subtrees are not equal. It is possible that $s^{i}\neq s^{i+1}$. Note that here the superscripts are relative to $b_{\ell}$ as opposed to Fig. 1 and Fig. 6.
As mentioned in the beginning of Section 2.3.1, only for simplicity we consider a symmetric setting in terms of sampled actions and the observations, but the approach is applicable without any limitations to any given belief tree.
@@ 4.1 Resimplification strategy: PT-SL
We now discuss how the simplification level is updated recursively from the simplification level of pertaining reward bounds, and revisit the process to calculate the optimal policy and the corresponding bounds. For some belief node $b_{\ell}$ in the belief tree, consider the bounds $\overline{\hat{Q}}^{s^{j}}(b_{\ell},\{a^{j}_{\ell},\pi^{\ast}_{(\ell+1)+}\})$ and $\underline{\hat{Q}}^{s^{j}}(b_{\ell},\{a^{j}_{\ell},\pi^{\ast}_{(\ell+1)+}\})$ from (45) for different actions $a^{j}_{\ell}\in\mathcal{A}$, that partially overlap and therefore could not be pruned. Each policy tree corresponding to action $a^{j}_{\ell}$ can generally have its own simplification level $s^{j}$. We now iteratively increase the simplification level by $1$. This can be done for each of the branches, if $s^{j}$ is identical for all branches, or only for the branch with the coarsest simplification level.
@@
For the former case, i.e. $s^{i}=s^{j}$, we now have to adapt the simplification level of a child tree $i$ to $s^{j}+1$ by calculating the bounds $\overline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$. Here, our key insight is that, instead of calculating these bounds from scratch, we can re-use calculations between different simplification levels, in this case, from level $s^{i}$. As the bounds from that level are available, we can identify only the incremental part that is “missing” to get from simplification level $s^{i}$ to $s^{i}+1$, and update the existing bounds $\overline{\hat{V}}^{s^{i}}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$ to recover $\overline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+}),\underline{\hat{V}}^{s^{i}+1}(b^{i}_{\ell+1},\pi^{\ast}_{(\ell+1)+})$ exactly. The same argument applies also for bounds over momentary rewards. In Section 6.2.3 we apply this approach to a specific simplification and reward function.